OpenStudy (anonymous):

O2(g) effuses at a rate that is _____ times that of Br2(g) under the same conditions. rate O2/rate Br2=

OpenStudy (aaronq):

Use the molar masses of the diatomic molecules in: \(\sf \dfrac{rate~_{O_2}}{rate~_{Br_2}}=\sqrt{\dfrac{Molar~mass~_{Br_2}}{Molar~mass~_{O_2}}}\)